A tank contains 78.0 g of N2 and 42.0 g of Ne at a total pressure of 3.75 atm and a temperature of 50.0°C. Calculate the following quantities.

(a) moles of N2

(b) moles of Ne

(c) the partial pressure of N2

(d) the partial pressure of Ne





1
Expert's answer
2021-10-31T00:04:21-0400

Solution:

The molar mass of N2 is 28.0134 g/mol

The molar mass of Ne is 20.1797 g/mol


Calculate the moles of each gas:

(a): (78.0 g N2) × (1 mol N2 / 28.0134 g N2) = 2.7844 mol N2 = 2.78 mol N2

Moles of N2 = 2.78 mol

(b): (42.0 g Ne) × (1 mol Ne / 20.1797 g Ne) = 2.0813 mol Ne = 2.08 mol Ne

Moles of Ne = 2.08 mol


Dalton's law of partial pressures can be used.

Dalton's law can be expressed using the mole fraction of a gas, x: Pgas1 = xgas1 × PTotal

The mole fraction (x) of a gas is the number of moles of that gas divided by the total moles of gas in the mixture:

xgas1 = mole fraction of gas 1 = (Moles of gas 1) / (Total moles of gas)


Total moles of gas = Moles of N2 + Moles of Ne = 2.78 mol + 2.08 mol = 4.86 mol


Calculate the mole fraction of each gas:

xN2 = (Moles of N2) / (Total moles of gas) = (2.78 mol) / (4.86 mol) = 0.572

xN2 = 0.572

xNe = (Moles of Ne) / (Total moles of gas) = (2.08 mol) / (4.86 mol) = 0.428

xNe = 0.428


Calculate the partial pressure of each gas:

(c): PN2 = xN2 × PTotal = 0.572 × 3.75 atm = 2.145 atm

The partial pressure of N2 = 2.145 atm

(d): PNe = xNe × PTotal = 0.428 × 3.75 atm =1.605 atm

The partial pressure of Ne = 1.605 atm


Answers:

(a): Moles of N2 is 2.78 mol

(b): Moles of Ne is 2.08 mol

(c): The partial pressure of N2 is 2.145 atm

(d): The partial pressure of Ne is 1.605 atm